A body projected vertically up with velocity root of 4gh stikes a glass plate at a height h above the ground from which it is projected.After breaking the glass it loses half of its velocity.Then it can go up to a further height of

Grade:11

1 Answers

Anil Pannikar AskiitiansExpert-IITB
13 years ago

Dear Chakrala,

 

v = sqrt(4gh) - 10t

h = sqrt(4gh)t - 5t2

solve for v and t,

then apply , 0 - v2 = - 2gH    here H = final attained height

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Anil Pannikar

IIT Bombay